# EVEv2: Improved Baselines for Encoder-Free Vision-Language Models

### v1

@@ 1 Introduction
With the recent rapid advancements in both large language models (LLMs) [56, 74, 89, 9, 8] and large vision models (LVMs) [20, 10, 22, 65, 18], vision-language models (VLMs) [15, 4, 52, 90, 67, 2] have made remarkable strides, demonstrating impressive capabilities in multi-modal understanding and reasoning applications. As illustrated in Figure 1(1), typical practice adopts pre-trained vision encoders to extract visual semantics, which are then translated into the text embedding space as Foreign language inputs for subsequent LLMs (e.g., BLIP [41] and LLaVA [48]). In contrast, another representative branch transforms visual features from vision encoders’ last layer across each layer of LLMs through cross-attention modules, like Flamingo [1] and LLaMA-3.2V [69]. Thanks to well-aligned representations across modalities through large-scale contrastive learning [58, 95, 63], these studies can achieve promising performance and strong multi-modality capability after joint training. However, the inductive biases during visual pre-training, e.g., image resolution, aspect ratio, and semantic priors, limit the flexibility and applicability of the visual learning in diverse real-world scenarios [87, 72, 71].
Unlike compositional VLMs, Fuyu [6] takes the early step to explore monolithic VLMs at different scales, eliminating the requirements for pre-trained visual encoders, while EVE [19] first pioneers a transparent, efficient, and practical path for advancing encoder-free VLM direction. Among them, PaliGemma [7] constructs an encoder-free VLM that demonstrates strong scaling efficiency while progressively approaching its encoder-based counterpart. With extensive training data and computing resources, Mono-InternVL [54] narrows the gap and matches the performance of Intern-VL1.5 [15] starting with the same LLM capabilities.
@@
Figure 2: Preliminary analyses across various VLMs’ scaling efficiency during pre-training or fine-tuning (More details in Appendix A). Notably, VE / DT / EVE apply varying image downsampling rates (142 / 82 / 322). For fairness, we choose slightly different resolutions that yield relatively balanced token counts of 576 / 1024 / 625 tokens per image. Besides, we quantify weight changes between LLMs and VLMs by averaging absolute value variation within specific layer number or type. We report accuracy on GQA [30], SEED [26], TextVQA [60], and SQA [53] to examine VLMs’ capabilities across general in-domain, open-domain, OCR-related, and text-related knowledge tasks.
@@ 2 Related Work
Encoder-free VLMs. Another visual processing strategy is discrete visual tokenization [75, 21, 61], which is widely used in various multi-modal understanding and generation approaches [66, 96, 81, 80]. However, the discretization inevitably results in lossy visual information and weakens in extracting semantic contents, which in turn hinders fine-grained visual understanding and reasoning [85, 82]. Therefore, recent studies [83, 42, 57, 86] introduce semantic constraints in the visual tokenizer for both high-level semantic and fine-grained visual representations. Compared to their highly-compressed features in the discrete-value space, encoder-free VLMs [6, 19, 13] have emerged as a promising architecture for lossless visual encoding, efficient data scaling, and end-to-end multimodal processing. Specially, EVE [19] pioneers an efficient and transparent path for monolithic VLMs, with its data-scaling efficiency preliminarily validated by PaliGemma [7]. Impressively, Mono-InternVL [54] bridges the performance gap with its modular counterpart [15] of the same LLM capacity, using adequate data. We emphasize that limited by current training data and device resources, EVEv2.0 does not aim for state-of-the-art performance but instead focuses on revealing the most efficient route for encoder-free VLMs from scratch.
@@ 3.1 Preliminary
Pioneer Experiments. Building on [47, 80, 19], we conduct two pilot experiments in Figure 2. Exp.(i): we adopt Vicuna-7B [16] via the vision encoder (CLIP-ViT-L-336px [58]), discrete tokenizer (Mo-VQGAN [99]), or one single transformer block [76] from scratch, dubbed VE, DT, or EVEv1.0 (w/o visual supervision). We first train the projector, vision vocabulary, or vision block using EVE-cap-16M [19] caption data, followed by EVE-cap-33M [19] to update only vision encoder for VE (work best), or extra LLM weights for DT and EVEv1.0, respectively. Finally, we update the overall backbone except Mo-VQGAN during instruction-tuning. Exp.(ii): we employ stronger Qwen2-7B [3] as the LLM for VE and EVEv1.0, compared with EVEv1.2 in Figure 1(2b) with patch embedding layer. Using 29M image-text annotations, we initially train the projector for VE or visual layers for EVE, with further updates to EVE’s LLM layers via extra 48M data. Subsequently, we import varying-scale instruction data [47, 19, 38] to fine-tune the entire network.
@@
Finding (2): Challenges towards multimodal interference and smooth transition. Exp.(ii) shows that compared with EVEv1.2 using the same data and stage, EVEv1.0 struggles to approach VE in text-related knowledge (SQA-IMG) due to catastrophic linguistic forgetting in LLMs. While mixing text-only and multi-modal data slightly alleviates this issue, it slows down the development of multi-modal understanding in Figure 7. Hence, we explore potential architectures and training strategies targeting LLM adaptation and multimodal interference. EVEv1.2 with re-parameterization design for each feed-forward weight helps smooth the transition from LLMs to VLMs, while EVEv1.5 with MoE design helps decouple vision-language encoding heterogeneity. However, after comparing LLMs and VLMs, we observe that VE requires only minor LLM adjustments to achieve robust results, whereas EVE necessitates extensive updates for similar capabilities in Figure 2. Besides, Layer normalization stands out as the most impacted module, indicating that EVEv1.2 and EVEv1.5 face challenges in efficiently constructing visual representations and achieving multimodal alignment when LLM weights are fixed in Figure 5. This dependency on the pre-training paradigm of LLM limits the VLMs’ full potential for learning visual perception from scratch.
@@ 3.3 Training Procedure
DenseFusion++. Developing strong visual perception requires high-quality, highly-detailed image-text data. Hence, we propose an efficient and low-budget engine based on LLaVA-1.6 (7B) [49] for scaling up synthetic data. Specifically, it leverages multiple vision experts (Tag, Detection, OCR, …) and learns GPT-4V’s fusion strategy, greatly facilitating data-scaling efficiency for native VLMs with high-quality annotations, not just distilling LLaVA. In Figure 6, we empirically demonstrate the superior training efficiency achieved by using intensive annotations from DenseFusion++ during pre-training. It consistently outperforms raw web-sourced captions or those generated by lower-quality caption engines. We will release the caption engine code and weight for better interpretation and further exploration.
@@ 2.2 Infinity-MM-GeneralQA [27] 15M 15M
Supervised Fine-tuning. During the SFT stage, we further enhance EVE’s ability to understand complex linguistic instructions and multifarious dialogue patterns, which are crucial for real-world applications. Here, we optimize the overall network architecture on a diverse set of high-quality, multi-source instruction datasets, namely EVE-sft-7M, including LLaVA-onevision [38] and partial Infinity-MM-instruct [27]. Notably, Stages 2.2 and 3 can be merged if large, balanced, and high-quality SFT data is available. We separate them to handle diverse but uneven (Stage 2.2) and balanced (Stage 3) data to achieve consistent performance.
@@ 4.2 Main Results
From Table 2, EVEv2.0 surpasses the encoder-free counterparts, e.g. Fuyu [6], EVE [19], SOLO [13], etc. across various vision-language benchmarks. Note that, due to training data and device resource limitations, we have not yet trained smaller models under 2B parameters in parallel and, as such, cannot provide a more direct comparison with Mono-InternVL [54]. However, we emphasize that our Divide-and-Conquer architecture demonstrates superior data-scaling efficiency compared to conventional Mixture-of-Experts designs, as shown in Figure 5. Preliminary experiments in Figure 2 reveal that simply decoupling the feed-forward module is insufficient to resolve vision-language conflicts and module compatibility inside one single unified network.
@@
Figure 6: Evaluation results of different data sources and caption engines. We utilize EVEv1.0 based on Vicuna-7B [16] as the baseline. Here ”*-raw“, ”*-cap“, or ”*-recap“ denote noisy web image captions, the samples annotated by both LLaVA-1.5 (13B) and Emu2 (17B), or modified DenseFusion++ (7B), respectively. Note that ”L.O.S.“ represents the mixture of LAION [59], OpenImages [35], and SAM [34].
Besides, EVEv2.0 displays superior performance against the VLMs using discrete tokenizers, i.e. Chameleon [66] and Emu3 [80], despite being trained on significantly fewer data or utilizing fewer visual tokens. This further validates the efficiency and effectiveness of encoder-free VLMs with lossless visual encoding mode, even using a lightweight patch embedding layer from scratch. Notably, EVEv2.0 competes with popular and mainstream encoder-based VLMs, e.g. LLaVA-1.6 [49] and Cambrian [71]. We argue that the performance gap between EVEv2.0 and advanced modular VLMs primarily arises from the significant discrepancy in data magnitude, restricting encoder-free VLMs from constructing a generalizable vision perception capability and handling more complicated visual perception scenarios.
@@ 4.3 Ablation Studies
Divide-and-Conquer (DaC) design outperforms the re-parameterization (ReP) and mixture-of-experts (MoE). As shown in Figure 5, (1) the training process of EVEv1.0 is the slowest, and training only the patch embedding layer proves insufficient, resulting in minimal performance gains. (2) EVEv1.2 (Rep) shows a rapid loss decrease, which can be attributed to the gradual transfer of LLMs into the initial VLMs by updating the pre-trained LLM weights. However, this approach leads to a noticeable performance drop on the SQA-IMG task requiring abundant text-related knowledge. (3) In contrast, EVEv1.5 (MoE) only updates visual parameters inside frozen LLMs to effectively mitigate catastrophic forgetting issues during pre-training. However, solely decoupling FFN modules restricts distinct feature distributions across modalities, resulting in less-efficient improvements in visual perception and multi-modality alignment. (4) With prior validation support, EVEv2.0 (DaC) achieves optimal improvements across all multi-modal benchmarks, highlighting its superior data-scaling efficiency during large-scale pre-training. This success can be attributed to its modality-wise sparsity, which effectively preserves linguistic knowledge while providing greater flexibility for visual learning. This philosophy is further evidenced by the loss curve in Figure 5 with faster convergence and better training stability than EVEv1.5 during pre-training. Notably, their Avg. accuracy gap rises from 0.8% to 1.4% as the training data grows from 8M to 24M, a trend likely to hold for other model sizes and data sources. Besides, only decoupling LayerNorm yields the Avg. accuracy of 48.8% vs. 51.6% for EVEv2.0 using 8M data, necessitating the complete decomposition.
@@
Figure 9: Evaluation results across progressive training procedure in Figure 9. We adopt standard EVEv2.0 based on Qwen-2.5 [70].
@@ 6 Conclusion
In this paper, we present EVEv2.0, a carefully designed and transparent encoder-free architecture designed for vision-and-language understanding and reasoning. Rather than focusing solely on state-of-the-art performance, we systematically analyze and identify the most efficient approach for building visual perception from scratch. To address interference and compatibility challenges between vision and language, we fully disentangle the model components and introduce modality-wise sparsity within a unified decoder-only backbone. Besides, we establish an efficient pathway for optimizing data-scaling efficiency in monolithic VLM research through a modified caption engine and a carefully designed training recipe. Using only 100M publicly available data, EVEv2.0 outperforms existing encoder-free models and steadily approaches the performance of encoder-based counterparts of similar capacity across a range of vision-language benchmarks. This provides valuable insights for developing scalable, native VLMs for the next generation.
@@ Appendix A Experiment Details
All experiment details in the main body are listed in EVEv2: Improved Baselines for Encoder-Free Vision-Language Models. ”EVE-cap-16M“ denotes the data mixture of LAION [59], OpenImages [35], and SAM [34] annotated by LLaVA-1.5 (13B) and Emu2 (17B). ”EVE-recap-16M“ denotes the data mixture of Datacomp [24], LAION [59], OpenImages [35], and SAM [34] annotated by DenseFusion++ (7B).
For Exp.(i) in Figure 2, we first use EVE-cap-16M in Stage 1 to train the projector, vision vocabulary embeddings, and lightweight vision block for the vision encoder (VE), the discrete tokenizer (DT), and EVEv1.0. In Stage 2, we use EVE-cap-33M to train only the vision encoder and projector for VE, as unfreezing LLM weights at this stage leads to performance collapse [19]. For DT and EVEv1.0, we unfreeze all model parameters. In Stage 3, we train all model weights across all models. Finally, we quantify weight changes between Vicuna-7B [16] and VE/EVEv1.0 to analyze the architectural differences. To ensure fairness, we remove the original vision encoder supervision in EVEv1.0.

### v2

@@ 1 Introduction
With the recent rapid advancements in the large language models (LLMs) [59, 77, 92, 9, 8] and large vision models (LVMs) [22, 24, 68, 18, 44], vision-language models (VLMs) [4, 55, 93, 2, 20, 21] have made remarkable strides, demonstrating impressive capabilities in multi-modal understanding and reasoning applications. As illustrated in Figure 1(1), typical practice adopts pre-trained vision encoders to extract visual semantics, which are then translated into the text embedding space as “Foreign” language inputs for subsequent LLMs (e.g., BLIP [43] and LLaVA [51]). In contrast, another representative branch transforms visual features from vision encoders’ last layer across each layer of LLMs through cross-attention modules, like Flamingo [1] and LLaMA-3.2V [72]. Thanks to well-aligned representations across modalities through large-scale contrastive learning [61, 98, 66], these studies can achieve promising performance and strong multi-modality capability after joint training. However, the inductive biases during visual pre-training, e.g., image resolution, aspect ratio, and semantic priors, limit the flexibility and applicability of the visual learning in diverse real-world scenarios [90, 75, 74].
Unlike them, Fuyu [6] takes an early step to explore monolithic VLMs via removing pre-trained visual encoders, while EVE [19] first pioneers a transparent, efficient, and practical path for advancing the encoder-free VLM direction. The subsequent PaliGemma [7] constructs an encoder-free VLM that shows strong scaling efficiency while progressively approaching its encoder-based counterpart. With extensive training data and computing resources, Mono-InternVL [57] narrows the gap and matches the performance of Intern-VL1.5 [15], starting with the same LLM capabilities.
@@
Figure 2: Preliminary scaling efficiency analyses during pre-training or fine-tuning across various VLMs. (More details in the Appendix). Notably, VE / DT / EVE apply varying image downsampling rates (142 / 82 / 322). For fairness, we choose different resolutions that yield relatively balanced token counts of 576 / 1024 / 625 tokens per image. Besides, we quantify weight changes between LLMs and VLMs by averaging absolute value variation within specific layer number or type. We report GQA [32], SEED [28], and TextVQA [63] for in-domain, open-domain, and OCR-related validation. Note that SQA [56] involves text-related knowledge tasks susceptible to LLM’s forgetting issue.
@@ 2 Related Work
Encoder-free VLMs. Another visual processing strategy is discrete visual tokenization [78, 23, 64], which translates the image into a sequence of discrete tokens and is widely used in various multi-modal understanding and generation approaches [69, 99, 84, 83]. However, the discretization inevitably results in lossy visual information and weakens in extracting semantic contents, which in turn hinders fine-grained visual understanding and reasoning [88, 85]. Therefore, recent studies [86, 45, 60, 89] introduce semantic constraints in the visual tokenizer for both high-level semantic and fine-grained visual representations. Compared to their highly-compressed features in the discrete-value space, encoder-free VLMs [6, 19, 13] have emerged as a promising architecture for lossless visual encoding, efficient data scaling, and end-to-end multimodal processing. Specially, EVE [19] pioneers an efficient and transparent path for monolithic VLMs, with its data-scaling efficiency preliminarily validated by PaliGemma [7]. Impressively, Mono-InternVL [57] bridges the performance gap with its modular counterpart [15] of the same LLM capacity, using adequate data. We emphasize that limited by current training data and device resources, EVEv2.0 does not aim for state-of-the-art performance but instead focuses on revealing the most efficient route for encoder-free VLMs from scratch.
@@ 3.1 Preliminary
Settings. Building on [50, 83, 19], we conduct two pilot experiments in Figure 2. Exp.(i): we combine Vicuna-7B [16] with vision encoder (VE: CLIP-ViT-L-336px [61]), discrete tokenizer (DT: Mo-VQGAN [102]), or one single transformer block [79] (EVEv1.0: w/o VE supervision). We first train the projector, vision vocabulary, or vision block using EVE-cap-16M [19] caption data, followed by EVE-cap-33M [19] to update only vision encoder for VE (work best), or extra LLM weights for DT and EVEv1.0. Finally, we update all layers except for discrete tokenizers during fine-tuning. Exp.(ii): we use stronger Qwen2-7B [3] as the LLM for VE, EVEv1.0, and EVEv1.2. Using 29M image-text pairs, we initially train the projector for VE or all visual-related layers for EVE, with further updates to EVE’s LLM layers via an extra 48M data. We then import varying-scale instruction data [50, 19, 40] to train the entire network.
@@
Finding (2): Challenges towards multimodal interference and smooth transition. While mixing text-only and multi-modal data can alleviate this issue, it slows down the development of multi-modal understanding [55]. Exp.(ii) shows that compared to EVEv1.0, EVEv1.2 approaches VE in knowledge-related SQA-IMG metric, slightly mitigating catastrophic linguistic forgetting in LLMs. Hence, we explore potential architectures and training strategies targeting LLM adaptation and multimodal interference. EVEv1.2 with re-parameterization design for each feed-forward weight helps smooth the transition from LLMs to VLMs, while EVEv1.5 with MoE design helps decouple vision-language encoding heterogeneity. However, after comparing LLMs and VLMs, we observe that VE only requires minor adjustments to the LLM, whereas EVE necessitates extensive updates for similar capabilities in Figure 2. Besides, Layer normalization emerges as the most impacted module, indicating that EVEv1.2 and EVEv1.5 struggle to efficiently construct visual representations and multi-modal alignments when attention and normalization layers in LLM are frozen. This dependency on LLM’s pre-training paradigm limits the full potential for learning visual perception from scratch.
@@ 3.3 Training Procedure
DenseFusion++. Developing strong visual perception requires high-quality, hyper-detailed image-text training data. Hence, we propose an efficient and low-budget engine based on LLaVA-1.6 (7B) [52] for scaling up synthetic data. Specifically, it leverages multiple vision experts (e.g., Tag, Detection, OCR, …) and learns the fusion strategy from GPT-4V, greatly facilitating data-scaling efficiency for native VLMs with high-quality annotations, not just distilling LLaVA. We empirically validate the superior training efficiency achieved by using intensive annotations from DenseFusion++ during pre-training. It consistently outperforms original web-sourced captions or those generated by lower-quality caption engines. We will release the code and weight of the caption engine for better interpretation and further exploration.
@@ 2.2 Infinity-MM-GeneralQA [29] 15M 15M
Supervised Fine-tuning. During the SFT stage, we further enhance EVE’s ability to understand complex linguistic instructions and multifarious dialogue patterns, which are crucial for real-world applications. Here, we optimize the overall network layers on a diverse set of high-quality, multi-source instruction datasets, namely EVE-sft-7M, including LLaVA-onevision [40] and partial Infinity-MM-instruct [29]. Note that Stages 2.2 and 3 can be merged if large, balanced, and high-quality instruction data is available. We separate them to handle diverse but uneven (Stage 2.2) and balanced (Stage 3) data to achieve consistent performance.
@@ 4.2 Main Results
In Table 2, EVEv2.0 surpasses encoder-free counterparts, e.g. Fuyu [6], EVE [19], SOLO [13], etc. across various vision-language benchmarks. Besides, Mono-InternVL [57], despite its smaller model size, achieves strong performance by leveraging 13x more pretraining data and large-scale, high-quality SFT data. Notably, preliminary experiments in Figure 2 confirm that merely decoupling the feed-forward module is insufficient to address vision-language conflicts and module compatibility within a single unified network. Furthermore, we clarify and validate that our Divide-and-Conquer architecture can achieve better data-scaling efficiency over the Mixture-of-Experts approach employed in Mono-InternVL [57], as demonstrated in Figure 5.
@@
Figure 6: Evaluation results of different data sources and caption engines. We utilize EVEv1.0 based on Vicuna-7B [16] as the baseline. Here “*-raw”, “*-cap”, or “*-recap” denote noisy web image captions, the samples annotated by both LLaVA-1.5 (13B) and Emu2 (17B), or modified DenseFusion++ (7B), respectively. Note that “L.O.S.” represents the mixture of LAION [62], OpenImages [37], and SAM [36].
Besides, EVEv2.0 displays superior performance against the VLMs using discrete tokenizers, i.e. Chameleon [69] and Emu3 [83], despite being trained on significantly fewer data or utilizing fewer visual tokens. This further validates the efficiency and effectiveness of encoder-free VLMs with lossless visual encoding mode, even using a lightweight patch embedding layer from scratch. Notably, EVEv2.0 competes with popular and mainstream encoder-based VLMs, e.g. LLaVA-1.6 [52] and Cambrian [74]. The performance gap with advanced modular VLMs stems largely from the significant discrepancy in in data scale. EVEv2.0 addresses this through modality-aware decomposition with high-quality annotations, offering an efficient and practical pathway for native VLMs to compete with encoder-based models.
@@ 4.3 Ablation Studies
Divide-and-conquer (DaC) design outperforms the re-parameterization (ReP) and mixture-of-experts (MoE). As shown in Figure 5, (1) the training process of EVEv1.0 is the slowest, and training only the patch embedding layer proves insufficient, resulting in minimal performance gains. (2) EVEv1.2 (ReP) shows a rapid loss decrease, which can be attributed to the gradual transfer of LLMs into the initial VLMs by updating the pre-trained LLM weights. However, this approach leads to a noticeable performance drop on the SQA-IMG task requiring abundant text-related knowledge. (3) In contrast, EVEv1.5 (MoE) only updates visual parameters inside frozen LLMs to effectively mitigate catastrophic forgetting issues during pre-training. However, solely decoupling FFN modules restricts distinct feature distributions across modalities, resulting in less-efficient improvements in visual perception and multi-modality alignment. (4) With prior validation support, EVEv2.0 (DaC) achieves optimal improvements across all multi-modal benchmarks, highlighting its superior data-scaling efficiency during large-scale pre-training. This success can be attributed to its modality-wise sparsity, which effectively preserves linguistic knowledge while providing greater flexibility for visual learning. This philosophy is further evidenced by the loss curve in Figure 5 with faster convergence and better training stability than EVEv1.5 during pre-training. (5) Notably, their Avg. accuracy gap rises from 0.8% to 1.4% as the training data grows from 8M to 24M, a trend likely to hold for other model sizes and data sources. Besides, only decoupling LayerNorm yields the Avg. accuracy of 48.8% vs. 51.6% for EVEv2.0 using 8M data, necessitating the complete decomposition.
@@
Figure 7: Evaluation results of different methods. We report results on GQA, SEED-ALL, TextQA, SQA-IMG, MMB, and MMVet.
@@
EVEv2.0 achieves notable performance, rivaling the modular backbone equipped with strong visual encoders. In Figure 7, we utilize Qwen-2.5 [73] as the LLM with EVE-recap-48M for pre-training and EVE-sft-7M for fine-tuning. (1) EVEv2.0, with its modality-specific components and enhanced caption engine, effectively approaches the performance of encoder-based LLaVA-1.5 [50]. It surpasses the prototype EVE [19], which relies on visual supervision and a unified dense framework, as well as Mono-InternVL [57], which suffers from insufficient architecture decomposition. (2) The DenseFusion++ delivers consistent and significant gains across various types of VLMs, leveraging diverse visual experts to boost generalist VLMs’ perception and data-scaling efficiency for native VLM construction. These findings further clarify the philosophy behind EVEv2.0.
@@
Figure 9: Evaluation results across progressive training procedures. We adopt standard EVEv2.0 based on Qwen-2.5 [73] as the LLM.
@@ 6 Conclusion
In this paper, we propose EVEv2.0 with a transparent and powerful encoder-free architecture for vision-and-language understanding. Rather than focusing solely on state-of-the-art results, we systematically analyze the most efficient strategies for building visual perception from scratch. To address vision-language interference, we fully disentangle the model components and introduce modality-wise sparsity within a unified decoder-only backbone. Besides, we establish an efficient pathway for optimizing data-scaling efficiency in monolithic VLM research through a modified caption engine and a meticulous training recipe. Utilizing only 100M public data, EVEv2.0 outperforms existing encoder-free counterparts and steadily approaches mainstream encoder-based approaches of similar model capacity across various vision-language benchmarks. This provides valuable insights for developing scalable, native VLMs for the next generation.
Acknowledgements. The paper is supported in part by the National Natural Science Foundation of China under grant No.62441231, 62293542, Liao Ning Province Science and Technology Plan No.2023JH26/10200016, Dalian City Science and Technology Innovation Fund No.2023JJ11CG001, and Natural Science Foundation of Zhejiang Province under Grant No.LD25F020001, Ningbo Key R&D project under Grant No.2025Z039, and National Key R&D Program of China (2022ZD0116302).
@@ Appendix A Experiment Details
All experiment details in the main body are listed in EVEv2: Improved Baselines for Encoder-Free Vision-Language Models. “EVE-cap-16M” denotes the data mixture of LAION [62], OpenImages [37], and SAM [36] annotated by LLaVA-1.5 (13B) and Emu2 (17B). “EVE-recap-16M” denotes the data mixture of Datacomp [26], LAION [62], OpenImages [37], and SAM [36] annotated by DenseFusion++ (7B).
For Exp.(i), we first use EVE-cap-16M in Stage 1 to train the projector, vision vocabulary embeddings, and lightweight vision block for the vision encoder (VE), the discrete tokenizer (DT), and EVEv1.0. In Stage 2, we use EVE-cap-33M to train only the vision encoder and projector for VE, as unfreezing LLM weights at this stage leads to performance collapse [19]. For DT and EVEv1.0, we unfreeze all model parameters. In Stage 3, we train all model weights across all models. Finally, we quantify weight changes between Vicuna-7B [16] and VE / EVEv1.0 to analyze the architectural differences. To ensure fairness, we remove the original vision encoder supervision in EVEv1.0.
For Exp.(ii), all VLMs use stronger Qwen2-7B [92] and high-quality EVE-recap. In Stage 1, we train the projector for VE, patch embedding layer for EVEv1.0, patch embedding and extra vision layer inside the LLM for EVEv1.2. In Stage 2 and 3, we train all model weights for EVEv1.0-1.2. Note that we skip Stage 2 for VE as the baseline for comparison. Here, we compare the weights between Qwen2 and VE / EVEv1.2 trained by LLaVA-onevision [40].
Table 4: Dataset details in Stage 2.2, and 3 for fine-tuning EVEv2.0. Note that ***-FL denotes the filtered training dataset.
Stage Dataset #Data
2.2 Cambrian-FL [74], Infinity-Instruct-FL [29], 5M
LVIS-Instruct-FL [80], Sharegpt4v-FL [12],
ALLaVA-laion-FL [11], ALLaVA-vflan-FL [11],
LLaVA-Pretrain-FL [50], DocReason-FL [31],
DocDownstream-FL [31], DocStruct4M-FL [31].
3 LLaVA-onevision [40], Infinity-MM-Synthesis [29], 7.3M
Infinity-MM-Preference [29], Infinity-Instruct-FL [29],
DenseFusion [46], Cambrian-FL [74], Docmatix-FL [38],
LVIS-Instruct-FL [80], BLIP-OCR[17], LLaVA-mix [50].
Appendix B Dataset Details
From Table 4, we use Cambrian-FL [74], Infinity-Instruct-FL [29], LVIS-Instruct-FL [80], Sharegpt4v-FL [12], ALLaVA-laion-FL [11], ALLaVA-vflan-FL [11], LLaVA-Pretrain-FL [50], DocReason-FL [31], DocDownstream-FL [31], and DocStruct4M-FL [31] in Stage 2.2. In Stage 3, we adopt LLaVA-onevision [40] and Infinity-MM [29].
Appendix C Hyper-parameter Configurations